# In-Situ Fine-Tuning of Wildlife Models in IoT-Enabled Camera Traps for Efficient Adaptation

### v2

@@ Abstract.
Resource-constrained IoT devices increasingly rely on deep learning models for inference tasks in remote environments. However, these models experience significant accuracy drops due to domain shifts when encountering variations in lighting, weather, and seasonal conditions. While cloud-based retraining can address this issue, many IoT deployments operate with limited connectivity and energy constraints, making traditional fine-tuning approaches impractical. We explore this challenge through the lens of wildlife ecology, where camera traps must maintain accurate species classification across changing seasons, weather, and habitats without reliable connectivity. We introduce WildFiT, an autonomous in-situ adaptation framework that leverages the key insight that background scenes change more frequently than the visual characteristics of monitored species. WildFiT combines background-aware synthesis to generate training samples on-device with drift-aware fine-tuning that triggers model updates only when necessary to conserve resources. Through extensive evaluation on multiple camera trap deployments, we demonstrate that WildFiT significantly improves accuracy while greatly reducing adaptation overhead compared to traditional approaches.
@@ 1. Introduction
Second, drift-aware fine-tuning leverages these synthesized images to enable a powerful new approach to model adaptation. Rather than waiting to observe performance degradation on real animal appearances, WildFiT proactively evaluates model accuracy using synthesized images that reflect current environmental conditions and species distributions. This allows camera traps to anticipate and address potential accuracy drops before they impact wildlife monitoring, while carefully managing computational resources by triggering fine-tuning only when necessary.
We conducted extensive evaluations of WildFiT across three camera trap datasets and three platforms. Our results demonstrate that WildFiT’s background-aware data synthesis surpasses other computationally efficient approaches (yun2019cutmix, ; zhang2017mixup, ) by 7.3% and diffusion model-based synthesis (objectstitch, ; controlcom, ) by 3.0% in model accuracy, while being several orders of magnitude faster (60-150 milliseconds vs 300-600 seconds for a batch of 32 images). WildFiT’s drift-aware fine-tuning achieves Pareto optimality in terms of fine-tuning frequency and classification accuracy under drifts. It achieves up to 1.5% higher accuracy and requiring 50% fewer fine-tuning rounds compared to periodic fine-tuning. End-to-end results show that WildFiT significantly outperforms several domain adaptation approaches by 20-35%, even surpassing methods that utilize ground-truth labels of animals in the new location for fine-tuning. We also show that a single fine-tuning iteration on a batch of 32 images takes 2-68 seconds depending on the IoT platform, demonstrating its practicality for real-world deployments.
@@ 2. Background and Motivation
Camera Trap Applications. Camera traps are motion-triggered cameras that are widely used methods for ecological monitoring in remote, often inaccessible locations (bruce2024large, ). While cloud-based platforms can leverage powerful AI algorithms for species identification (ahumada2020wildlife, ), the high costs of field visits and data retrieval (estimated at \$2.15M for a typical monitoring program (smith2024man, )) have driven increasing interest in on-device processing. These devices leverage machine learning (ML) models for automated species identification, movement tracking, and rare species detection directly on the device (falzon2019classifyme, ). Processing data locally through on-device inference offers three key benefits: it eliminates the substantial costs and carbon footprint associated with frequent field visits (smith2024man, ), enables real-time responses to time-critical events like invasive species detection (ramsey2011quantifying, ), and ensures continuous monitoring even in areas with unreliable or non-existent network connectivity (glover2019camera, ; rastikerdar2024cactus, ).
@@
We see that both these shifts can have substantial impact on performance. A model trained on data from one location may perform poorly when deployed in a different environment with unfamiliar backgrounds or lighting conditions. In addition, the specific species of animals that are frequent in an area can vary across locations and seasons, further complicating the classification task.
@@ 3.1. Overview of WildFiT
Figure 3 illustrates the design of WildFiT. At its core, WildFiT runs a lightweight classification model that identifies animals captured by motion-triggered cameras. To maintain accuracy as environmental conditions change, WildFiT implements a Drift-Aware Fine-Tuning pipeline that continuously monitors and responds to domain shifts through three key components:
@@ 3.4. Drift Validation
However, achieving statistically significant validation requires evaluating the model on many synthesized images, which can be computationally prohibitive for resource-constrained devices. We address this challenge through a reuse mechanism that intelligently caches and reuses validation results. By maintaining statistics from previous validations, the system can efficiently estimate model performance on new domains while minimizing redundant computations.
@@
Through this dual-path approach, Drift Validation efficiently maintains model quality by responding to both environmental and animal distribution changes while minimizing computational overhead.
@@ 3.5. Model Fine-Tuning
To prevent catastrophic forgetting while keeping the solution practical for IoT deployments, we implement two simple yet effective techniques that are easy to deploy on IoT devices. First, we combine recent background images with historical background images to create a more diverse set of synthetic images. These images are stored in Background Images (target) illustrated in Figure 3. This ensures that the model retains knowledge of environmental variations over time, preventing it from becoming overly specialized to the most recent backgrounds. Second, we adjust the class distribution to prevent the model from over-fitting to recent class occurrences. Mathematically, let $p(c)$ represent the probability of an animal class $c$ appearing in the scene recently, and let $N_{c}$ denote the number of sampled images for class $c$ based on the class distribution. We sample animal classes using a softened distribution defined as $N^{s}_{c}=N_{c}+T$, where $T$ is a tunable temperature parameter that controls the degree of distribution smoothing. By default, we set $T=20$ throughout our evaluation, which we find to work effectively.
@@ 4. Evaluation
This section empirically evaluates the efficacy of WildFiT. We describe the experiment settings in § 4.1, and evaluate Background-Aware Data Synthesis and Drift-Aware Fine-Tuning in § 4.2-4.3. We perform end-to-end evaluations and report longitudinal performance of WildFiT in § 4.4-4.5. We finally report ablation studies in § 4.6.
@@ 4.1. Experiment Settings
Datasets. We utilize three datasets as shown in Table 2: two from the Serengeti Safari Camera Trap network (referred to as D1 and D2) (snapshotserengeti, ), and one from the Enonkishu Camera Trap dataset (referred to as D3) (snapshotenonkishu, ). These datasets are chosen since they have temporal data over relatively long durations and across several locations captured by trail cameras, hence they exhibit real-world temporal and spatial domain drifts. The Camera Trap datasets normally include a lot of species plus an empty class whose images show a scene with no species in it. Some species lack enough samples for both training and testing sets. Following prior practice (rastikerdar2024cactus, ), we selected the 18 most frequent species from the Serengeti dataset and 10 from the Enonkishu dataset for classification and included the empty class to result in 19 and 11 classes respectively. For simplicity, we focused on images containing only one type of animal. Table 2 summarizes the data statistics. “Train” refers to locations whose images are used for training, while “Test” represents those for evaluating domain shifts.
@@
WildFiT Training. The wildlife classifier is trained in PyTorch using a two-stage pipeline on the source domain data to ensure high model quality. It employs the EfficientNet-B0 model pre-trained on ImageNet (deng2009imagenet, ). In the first stage, the classification head is trained for 5 epochs with a learning rate (lr) of 1e-3 while keeping the backbone frozen. In the second stage, the entire network is trained for 30 epochs with an lr of 1e-5, using an early stopping criterion of 2 epochs. We use Adam optimizer (kingma2014adam, ) and a cross-entropy loss function.
@@
The fine-tuning process is triggered by two drift detection mechanisms. Class Distribution Drift Detection (CDD) is triggered after every $C=100$ new predictions. If the dominant class is the empty class, the class distribution is reset. For Drift Validation, we reserve 10% of objects per class exclusively for validation, while using the remaining data for fine-tuning. The model is fine-tuned only when the Drift Validation module detects a performance drop on synthesized data using the most recent backgrounds and current class distribution. For both BDD and CDD we set the p-value to 5% and $thr$ for the Drift Validation module is 0%. We utilized a 2-day window after the BDD triggers a drift and before the Drift Validation module is activated. This ensures that at least 2 days have passed since the last fine-tuning event before Drift Validation is triggered again, effectively controlling the frequency of Drift Validation for background drifts.
Platforms. We use a cluster of NVIDIA L40S 48GB GPUs for model training to prepare the initial wildlife classifier. While this offline training uses powerful GPUs, we implement WildFiT’s in-situ fine-tuning capabilities and evaluate its runtime performance on resource-constrained IoT devices, using EfficientNet-B0 on Raspberry Pi 5, Jetson Xavier NX and Orin AGX developer kits.
@@ Methods
Locs.  Ours  No FT $\Delta 1$  Mix Cut $\Delta 2$  Obj-St CC $\Delta 3$
@@
(1) Periodic fine-tuning (Periodic), where fine-tuning occurs at fixed intervals of N days. The fine-tuning process utilizes randomly sampled background images. To ensure a fair comparison, the number of randomly sampled backgrounds is matched to the number of backgrounds detected by BDD. A larger N results in less number of fine-tuning but risk higher accuracy drops under background drifts.
@@
Figure 6(a) illustrates the accuracy-efficiency tradeoffs achieved by different approaches. For the Periodic Fine-Tuning baselines, we vary the intervals between fine-tuning events, setting $N={2,4,6,8,10,14,21}$ days to achieve different trade-off points. In contrast, our approach adjusts the thresholds for acceptable accuracy degradation within the Drift Validation module, utilizing thresholds $thr$ from $-6\%$ to $0\%$ in $1\%$ increments to explore various trade-off scenarios.
@@
Figure 6(b) reports the trade-off between fine-tuning frequency and model accuracy for our proposed approach and the baselines under class distribution drift. For the Periodic Fine-Tuning baselines, the intervals between fine-tuning events are varied, with $D={4,7,10,14,18,21,28}$ days, to explore different trade-off points. For Triggered+CDD, the p-value in the p-test is varied, with $p={0.5\%,1\%,2\%,3\%,4\%,5\%}$. In contrast, our approach adjusts the thresholds for acceptable accuracy degradation within the Drift Validation module, utilizing thresholds $thr$ from $-6\%$ to $0\%$ in $1\%$ increments as before to explore various trade-off scenarios.
@@ 4.4. End-to-End Performance
(3) WildFiT:Syn: This approach removes the drift-aware fine-tuning for domain adaptation. We start by initializing our background repository with the first 80 samples and perform a single round of model fine-tuning. This baseline addresses the spatial domain shift only.
@@
Comparison with Pseduolabel: The Pseduolabel approach performs poorly in all datasets. The accuracy only reaches 55.4%, 58.3% and 39.3% for D1, D2 and D3, respectively. This underperformance is primarily due to the inaccuracy of the pseudo-labels as the starting model is only trained on the source domain, highlighting the importance of proactive fine-tuning rather than relying on reactive approaches.
@@
RPi-5 Xavier NX Orin AGX
WildFiT:Syn (CPU) 61.6 ms 149.7 ms 60.2 ms
CutMix:Syn (CPU) 69.9 ms 445.4 ms 36.7 ms
MixUp:Syn (CPU) 100.2 ms 446.5 ms 43.2 ms
Object-Stitch:Syn (GPU) - - 319.1 s
Control-Com:Syn (GPU) - - 613.2 s
WildFiT:Fine-Tuning 67.65 s 3.33 s 1.75 s
@@
Table 5. Latency for synthesis and Fine-Tuning (one iteration) on a batch of 32 images is compared across WildFiT and alternative approaches on three devices: Raspberry Pi 5 (8GB), Jetson Xavier NX (8GB), and Jetson Orin AGX (32GB). CPU-based approaches for image synthesis use #workers = 4.
@@
Runtime Performance Results. Table 5 reports the runtime performance of WildFiT on three platforms. To fine-tune the model on the Raspberry Pi 5 and Jetson Xavier NX with an effective batch size of 32 without OOM, we adopted a gradient accumulation technique that uses a batch size of 4 with 8 accumulation steps to update the weights. The synthesis latency is the time spent on the image synthesis and does not include the latency introduced by the dataloader.
@@
Overall, the results show that WildFiT has very low computational overhead and is highly practical for real-world use on resource-constrained platforms while allowing our wildlife classification system to continuously adapt to changing environments. In particular, WildFiT is extremely efficient in generating synthetic images, introducing minimal overhead for model fine-tuning. For example, on the Raspberry Pi 5, WildFiT spends 61.6 milliseconds for synthesizing a batch of 32 images, which is significantly faster than the 67.65 seconds required for training a single batch with EfficientNetB0. This high efficiency is consistent on more powerful devices, the Jetson Xavier NX and Jetson Orin AGX. Although data augmentation methods like CutMix and MixUp are comparably efficient, they result in worse model performance due to the lower quality of the synthesized data, as discussed in § 4.2.
In contrast, diffusion-based image composition techniques, despite utilizing GPUs, are exceedingly slow and cannot be executed on lower-tier devices like the Raspberry Pi 5 and Jetson Xavier NX. For instance, synthesizing a batch of 32 images using Control-Com, even on the Orin AGX, is thousands of times slower than WildFiT on lower-tier devices like the Raspberry Pi 5 and Xavier NX. Additionally, Control-Com image synthesis is 184$\times$ and 350$\times$ slower than training the model for a single iteration on the Xavier NX and Pi 5, respectively. This significant latency renders these techniques impractical on IoT platforms.
Results on Memory Requirements. The peak memory usage for the Jetson Orin AGX during fine-tuning with a batch size of 32, updating only the Fully Connected layers, biases, and Batch Normalization layers, is 15.1 GB for GPU memory and 5.5 GB for CPU memory. For the Raspberry Pi 5, which lacks a GPU, the peak CPU memory usage is 4 GB when fine-tuning with a batch size of 4 and using gradient accumulation to achieve an effective batch size of 32. For Xavier NX the peak GPU and CPU usage are 2 GB and 3.5 GB, respectively.

### v3

@@ Abstract.
Resource-constrained IoT devices increasingly rely on deep learning models, however, these models experience significant accuracy drops due to domain shifts when encountering variations in lighting, weather, and seasonal conditions. While cloud-based retraining can address this issue, many IoT deployments operate with limited connectivity and energy constraints, making traditional fine-tuning approaches impractical. We explore this challenge through the lens of wildlife ecology, where camera traps must maintain accurate species classification across changing seasons, weather, and habitats without reliable connectivity. We introduce WildFiT, an autonomous in-situ adaptation framework that leverages the key insight that background scenes change more frequently than the visual characteristics of monitored species. WildFiT combines background-aware synthesis to generate training samples on-device with drift-aware fine-tuning that triggers model updates only when necessary to conserve resources. Our background-aware synthesis surpasses efficient baselines by 7.3% and diffusion models by 3.0% while being orders of magnitude faster, our drift-aware fine-tuning achieves Pareto optimality with 50% fewer updates and 1.5% higher accuracy, and the end-to-end system outperforms domain adaptation approaches by 20-35% while consuming only 11.2 Wh over 37 days—enabling battery-powered deployment.
@@ 1. Introduction
Second, drift-aware fine-tuning leverages these synthesized images to enable a powerful new approach to model adaptation. Whenever a potential drift is detected, drift-aware fine-tuning first invokes a drift validation module to evaluate model accuracy using synthesized images that reflect current environmental conditions and species distributions. It triggers fine-tuning only when drift validation confirms accuracy drops. This allows camera traps to anticipate and address accuracy drops before they impact wildlife monitoring, while carefully managing computational resources by triggering fine-tuning only when necessary.
We conducted extensive evaluations across three camera trap datasets and four platforms. Our results show that
  • WildFiT’s background-aware data synthesis surpasses other computationally efficient approaches (yun2019cutmix, ; zhang2017mixup, ) by 7.3% and diffusion model-based synthesis (objectstitch, ; controlcom, ) by 3.0% in model accuracy, while being several orders of magnitude faster (10-150 ms vs 300-600 seconds for a batch of 32 images).
  • WildFiT’s drift-aware fine-tuning achieves Pareto optimality in terms of fine-tuning frequency and classification accuracy under drifts. It achieves up to 1.5% higher accuracy and requiring 50% fewer fine-tuning rounds compared to periodic fine-tuning.
  • End-to-end results show that WildFiT significantly outperforms several domain adaptation approaches by 20-35%, even surpassing methods that utilize ground-truth labels of animals in the new location for fine-tuning. We also show that a single fine-tuning iteration on a batch of 32 images takes 2-68 seconds depending on the IoT platform, demonstrating its practicality for real-world deployments.
[…]
  • The end-to-end WildFiT pipeline is highly energy-efficient and consumes merely 11.2 Wh for processing a real stream of camera trap images spanning 37 days—a consumption level supported by three alkaline AA batteries.
@@ 2. Background and Motivation
Camera Trap Applications. Camera traps are motion-triggered cameras that are widely used methods for ecological monitoring in remote, often inaccessible locations (bruce2024large, ). While cloud-based platforms can leverage powerful AI algorithms for species identification (ahumada2020wildlife, ), the high costs of field visits and data retrieval (estimated at \$2.15M for a typical monitoring program (smith2024man, )) have driven increasing interest in on-device processing. These devices leverage machine learning (ML) models for automated species identification, movement tracking, and rare species detection directly on the device (falzon2019classifyme, ). Processing data locally through on-device inference offers three key benefits: it eliminates the substantial costs and carbon footprint associated with frequent field visits (smith2024man, ), enables real-time responses to time-critical events like invasive species detection (ramsey2011quantifying, ), and ensures continuous monitoring even in areas with unreliable network connectivity (glover2019camera, ; rastikerdar2024cactus, ).
@@
We see that both these shifts can have a substantial impact on performance. A model trained on data from one location may perform poorly when deployed in a different environment with unfamiliar backgrounds or lighting conditions. In addition, the specific species of animals that are frequent in an area can vary across locations and seasons, further complicating the classification task.
@@ 3.1. Overview of WildFiT
Figure 3 illustrates the design of WildFiT. At its core, WildFiT runs a lightweight classification model that identifies animals captured by motion-triggered cameras on-device. To maintain accuracy as environmental conditions change, WildFiT implements a Drift-Aware Fine-Tuning pipeline that continuously monitors and responds to domain shifts through three key components:
@@ 3.4. Drift Validation
However, achieving statistically significant validation requires evaluating the model on many synthesized images, which can be computationally prohibitive for resource-constrained devices. We address this challenge through a reuse mechanism that intelligently caches and reuses validation results. By maintaining statistics from previous validations, the system can efficiently estimate model performance on new domains while minimizing redundancy.
@@
Through this approach, Drift Validation efficiently maintains model quality by responding to both environmental and animal distribution changes while minimizing computational overhead.
@@ 3.5. Model Fine-Tuning
To prevent catastrophic forgetting while keeping the solution practical for IoT deployments, we implement two simple yet effective techniques that are easy to deploy. First, we combine recent background images with historical background images to create a more diverse set of synthetic images. These images are stored in Background Images (target) illustrated in Figure 3. This ensures that the model retains knowledge of environmental variations over time, preventing it from becoming overly specialized to the most recent backgrounds. Second, we adjust the class distribution to prevent the model from over-fitting to recent class occurrences. Mathematically, let $p(c)$ represent the probability of an animal class $c$ appearing in the scene recently, and let $N_{c}$ denote the number of sampled images for class $c$ based on the class distribution. We sample animal classes using a softened distribution defined as $N^{s}_{c}=N_{c}+T$, where $T$ is a tunable temperature parameter that controls the degree of distribution smoothing. By default, we set $T=20$ throughout our evaluation, which we find to work effectively.
@@ 4. Evaluation
This section empirically evaluates the efficacy of WildFiT. We describe the experiment settings in § 4.1, and evaluate Background-Aware Data Synthesis and Drift-Aware Fine-Tuning in § 4.2-4.3. We perform end-to-end evaluations and report longitudinal performance of WildFiT in § 4.4-4.5, and ablation studies in § 4.6.
@@ 4.1. Experiment Settings
Datasets. We utilize three datasets as shown in Table 2: two from the Serengeti Safari Camera Trap network (referred to as D1 and D2) (snapshotserengeti, ), and one from the Enonkishu Camera Trap dataset (referred to as D3) (snapshotenonkishu, ). These datasets are chosen since they have temporal data over relatively long durations and across several locations captured by trail cameras, hence they exhibit real-world temporal and spatial domain drifts. The Camera Trap datasets normally include a lot of species plus an empty class whose images show a scene with no species in it. Since some species lack enough samples for both training and testing sets, we selected the 18 most frequent species from the Serengeti dataset and 10 from the Enonkishu dataset for classification and included the empty class to result in 19 and 11 classes respectively. For simplicity, we focused on images containing only one type of animal. Table 2 summarizes the data statistics. “Train” refers to locations whose images are used for training, while “Test” represents those for evaluating domain shifts.
@@
WildFiT Training. WildFiT is trained in PyTorch using a two-stage pipeline on the source domain data to ensure high model quality. It employs the EfficientNet-B0 model pre-trained on ImageNet (deng2009imagenet, ). In the first stage, the classification head is trained for 5 epochs with a learning rate (lr) of 1e-3 while keeping the backbone frozen. In the second stage, the entire network is trained for 30 epochs with an lr of 1e-5, using an early stopping criterion of 2 epochs. We use Adam optimizer (kingma2014adam, ) and a cross-entropy loss function.
@@
The fine-tuning process is triggered by two drift detection mechanisms. Class Distribution Drift Detection (CDD) is triggered after every $C=100$ new predictions. If the dominant class is the empty class, the class distribution is reset. For Drift Validation, we reserve 10% of objects per class exclusively for validation, while using the remaining data for fine-tuning. The model is fine-tuned only when the Drift Validation module detects a performance drop on synthesized data using the most recent backgrounds and current class distribution. For both BDD and CDD we set the p-value to 5% and thrthr for the Drift Validation module is 0%. We utilized a 2-day window after the BDD triggers a drift and before the Drift Validation module is activated. This ensures that at least 2 days have passed since the last fine-tuning event before Drift Validation is triggered again, effectively controlling the frequency of Drift Validation for background drifts.
Platforms. We use a cluster of NVIDIA L40S 48GB GPUs for model training to prepare the initial wildlife classifier. While this offline training uses powerful GPUs, we implement WildFiT’s in-situ fine-tuning capabilities and evaluate its runtime performance on resource-constrained IoT devices, using EfficientNet-B0 on Raspberry Pi 5, Jetson Orin Nano Super, Xavier NX, and Orin AGX developer kits.
Figure 6 shows the complete end-to-end wildlife camera system capable of capturing images and executing the WildFiT pipeline on the Jetson Orin Nano Super. In this example, images from the D2 dataset are replayed to demonstrate system functionality; however, the pipeline also supports direct input from the onboard camera. A demonstration video showcasing the fully operational end-to-end system is available at (wildfit2025demo, ).
Figure 6. Jetson Orin Nano Super with a camera, LCD display, and a power bank. The LCD display is optional.
@@ Methods
Locs.  Ours  No FT $\Delta 1$  Mix Cut $\Delta 2$  Obj-St $C$ $\Delta 3$
@@
(1) Periodic fine-tuning (Periodic), where fine-tuning occurs at fixed intervals of N days. The fine-tuning process utilizes randomly sampled background images. To ensure a fair comparison, the number of randomly sampled backgrounds is matched to the number of backgrounds detected by BDD. A larger N results in less fine-tuning but risk higher accuracy drops under background drifts.
@@
Figure 7(a) illustrates the accuracy-efficiency tradeoffs achieved by different approaches. For the Periodic Fine-Tuning baselines, we vary the intervals between fine-tuning events, setting N=2,4,6,8,10,14,21N={2,4,6,8,10,14,21} days to achieve different trade-off points. In contrast, our approach adjusts the thresholds for acceptable accuracy degradation within the Drift Validation module, utilizing thresholds thrthr from $6\%$ to $0\%$ in $1\%$ increments to explore various trade-off scenarios.
@@
Figure 7(b) reports the trade-off between fine-tuning frequency and model accuracy for our proposed approach and the baselines under class distribution drift. For the Periodic Fine-Tuning baselines, the intervals between fine-tuning events are varied, with D=4,7,10,14,18,21,28D={4,7,10,14,18,21,28} days, to explore different trade-off points. For Triggered+CDD, the p-value in the p-test is varied, with $p={0.5\%,1\%,2\%,3\%,4\%,5\%}$. In contrast, our approach adjusts the thresholds for acceptable accuracy degradation within the Drift Validation module, utilizing thresholds thrthr from $6\%$ to $0\%$ in $1\%$ increments as before to explore various trade-off scenarios.
@@ 4.4. End-to-End Performance
(3) WildFiT:Syn: This approach removes drift-aware fine-tuning for domain adaptation. We start by initializing our background repository with the first 80 samples and perform a single round of model fine-tuning. This baseline addresses the spatial domain shift only.
@@
Comparison with Pseduolabel: The Pseduolabel approach performs poorly in all datasets with accuracies 55.4%, 58.3% and 39.3% for D1, D2 and D3, respectively. This underperformance is primarily due to the inaccuracy of the pseudo-labels as the starting model is only trained on the source domain, highlighting the importance of proactive fine-tuning rather than relying on reactive approaches.
@@
RPi-5 O.S. Nano X. NX O. AGX
Synthesis Methods
WildFiT:Syn (CPU) 61.6 ms 10.2 ms 149.7 ms 60.2 ms
CutMix:Syn (CPU) 69.9 ms 6.9 ms 445.4 ms 36.7 ms
MixUp:Syn (CPU) 100.2 ms 9.2 ms 446.5 ms 43.2 ms
Object-Stitch:Syn (GPU) - - - 319.1 s
Control-Com:Syn (GPU) - - - 613.2 s
Fine-Tuning
WildFiT:Fine-Tuning 67.65 s 1.98 s 3.33 s 1.75 s
@@
Table 5. Latency for synthesis and Fine-Tuning (one iteration) on a batch of 32 images is compared across WildFiT and alternative approaches on four devices: Raspberry Pi 5 (8GB), Jetson Orin Nano Super (8GB), Jetson Xavier NX (8GB), and Jetson Orin AGX (32GB). CPU-based approaches for image synthesis use #workers = 4.
@@
Runtime Performance Results. Table 5 reports the runtime performance of WildFiT on four platforms representing a range of IoT deployment scenarios. We evaluated on the Raspberry Pi 5 for ultra-low-power deployments, the Jetson Orin Nano Super as a representative of current-generation edge AI devices that we expect to be widely deployed in future IoT systems, the Jetson Xavier NX for mid-range applications, and the Jetson Orin AGX for high-performance edge computing. To fine-tune the model on the Raspberry Pi 5, Jetson Orin Nano, and Jetson Xavier NX with an effective batch size of 32 without OOM, we adopted a gradient accumulation technique that uses a batch size of 4 with 8 accumulation steps to update the weights. The synthesis latency is the time spent on the image synthesis and does not include the latency introduced by the dataloader.
@@
Overall, the results show that WildFiT has very low computational overhead and is highly practical for real-world use on resource-constrained platforms while allowing our wildlife classification system to continuously adapt to changing environments. The performance on the Jetson Orin Nano Super is particularly promising, with synthesis taking only 10.2 milliseconds and fine-tuning requiring just 1.98 seconds per iteration, demonstrating excellent practicality for real-world IoT deployments. We see that WildFiT is extremely efficient in generating synthetic images, introducing minimal overhead for model fine-tuning. For example, on the Raspberry Pi 5, WildFiT spends 61.6 milliseconds for synthesizing a batch of 32 images, which is approximately three orders of magnitude faster than the 67.65 seconds required for training a single batch with EfficientNetB0. Similarly, on the Jetson Orin Nano Super, synthesis takes only 10.2 milliseconds, roughly three orders of magnitude less than the 1.98 seconds for fine-tuning. This high efficiency is consistent on other tested devices as well. Although data augmentation methods like CutMix and MixUp are comparably efficient for synthesis, they result in worse model performance due to the lower quality of the synthesized data, as discussed in § 4.2.
[…]
Results on Memory Requirements. Memory usage varies due to hardware constraints and batch size adaptation. The Orin AGX uses a batch size of 32 (15.1 GB GPU, 5.5 GB CPU), while other devices use a batch size of 4 with gradient accumulation for effective batch size of 32: Raspberry Pi 5 requires 4 GB CPU memory, Orin Nano Super uses 2 GB GPU and 2.5 GB CPU, and Xavier NX uses 2 GB GPU and 3.5 GB CPU memory. The difference in CPU memory usage between the Nano Super and Xavier NX reflects their different CPU architectures and cache efficiencies, with the Nano Super benefiting from a newer CPU generation.